Five final year science students from UNAM started from scratch to mix and filling of sanitisers in batches of 50 containers of 500ml in order to meet the unexpected spike in demand. Each intern is remunerated a rate of N$7.50 per hour of production. It takes 12 hours to manufacture the first batch and a total cumulative time of 27 hours to manufacture the 4th batch.

Required:

4.1Using the learning curve techniques, determine the rate of learning for the given scenario.

4.2Determine the time it take the team to complete the 8th batch.

4.3 The Ministry of Health and Social Services has placed a special order of 150 containers. What would be the incremental direct labor cost for the special order assuming normal production are 8 batches


1
Expert's answer
2022-03-17T09:39:24-0400

Computing the learning rate of the learning curve

1 batch "= 50" containers

It takes ; 12 hours

The 4th batch( 200 containers) takes 27 hours

200 units ;"Initial . time \u00d7 (2\u00d7lc)^3= 27"

"12\u00d7(2\u00d7lc)^3 = 27"

"96(lc)^3=27"

"lc = 0.655"


The time taken for the 8th batch will be;

"Initial.time \u00d7 (2\u00d70.655)^7"

"12\u00d7(2\u00d70.655)^7= 79.448" hours


150 containers will take

"12\u00d7(2\u00d70.655)^2= 20.59" hours

"79.448 - 20.59= 58.858" hours

"58.858\u00d77.5= N441.435"

This will be the incremental direct labour cost.